# Evaluating Visual and Cultural Interpretation: The K-Viscuit Benchmark with Human-VLM Collaboration

### v2

@@ 5.1 Evaluation beyond Multiple-choice VQA
To quantitatively analyze this aspect, we randomly sampled 80 questions from the Food category and evaluated Llava-v1.6-13B’s performance in a generative setting. The model was prompted with the instruction: "This is a question about Korean culture. Answer the given questions briefly and concisely." without access to the original multiple-choice options. Two native Korean evaluators assessed the generated responses using three categories: (1) Correct: The model’s response fully contained the original answer or was reasonably accurate to the question, (2) Hallucinated: The model confidently provided incorrect information, and (3) Abstained: The model’s response did not include a direct answer to the question but provided a relevant description of the image and did not attempt to generate an answer hastily (e.g., "The specific type of food and its cultural context would determine when it is commonly served").
The human analysis results are shown in Table 6. We observed that the VLM’s accuracy in the generative setup was relatively lower than in the multiple-choice setup (from 45.07% to 36.25%). This decrease in accuracy is likely because the task is inherently more difficult without predefined options. Interestingly, a significant proportion of responses (21.25%) were classified as abstained, indicating that the VLM recognized the need for additional information to answer confidently and refrained from providing a definitive but potentially incorrect response. This behavior suggests that providing the necessary external knowledge could enhance performance. Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
@@
Table 6: Human evaluation results of LLaVA-1.6-13B on Food category.
@@ 5.2 Augmenting VLMs with External Knowledge
As shown in Table 7, Providing retrieved documents can enhance model performance, as demonstrated by improved scores in the Retrieved setting compared to None. In cases where proprietary models did not benefit from certain retrieved documents, their performance still surpassed the baseline when given carefully curated Oracle documents. This suggests that retrieval-based augmentation has the potential to strengthen cultural knowledge in VQA tasks, provided that the quality and relevance of the selected documents are refined. We provide the details in the Appendix D.
@@
Table 7: Retrieval-augmented Generation Results.
@@ 6 Conclusion
In this work, we proposed a semi-automated framework for constructing culturally aware benchmarks for vision-language models through human-VLM collaboration, focusing on visual recognition and reasoning in multi-choice QA. We demonstrated its effectiveness by creating K-Viscuit, revealing a significant performance gap between proprietary and open-source VLMs in understanding Korean culture. Through detailed analyses and knowledge augmentation experiments, we established the impact of cultural understanding on VQA performance and extended our investigation to open-ended answer generation. This work highlights the importance of cultural diversity in model evaluation and paves the way for more inclusive VLMs.
@@
Our current framework requires a manual selection of images that match specified concepts, preventing fully automated dataset generation. These human efforts can be alleviated by multimodal retrieval modules to some extent. To this end, it should be predetermined whether current multimodal encoders can sufficiently understand culturally nuanced images and texts. Enhancing retrieval models to better understand and match cultural contexts remains our exciting future work. The manual verification of automatically generated questions also can be a considerable burden. Developing a quality estimation module for generated questions could assist in this process by reducing the workload on human annotators.
@@
Appendix A Dataset Construction Details
Guidelines to GPT-4-Turbo for Dataset Annotation
@@ Appendix A Dataset Construction Details
Table 9: A prompt of GPT-4-Turbo used in the annotation of Type 2 questions (i.e., visual reasoning) in the Architecture category.
Appendix B Dataset Analyses Details
B.1 Required Knowledge Analysis
@@
B.2 Human Evaluation Details
@@
Appendix C VLM Evaluation Details
@@
Appendix D Retrieval Methodology Details
@@
Appendix E Comparison of K-Viscuit with CVQA dataset

### v3

@@ 5.1 Evaluation beyond Multiple-choice VQA
Table 6: Generative QA Results. We manually evaluated the responses of LLaVA-1.6-13B on Food category.
To quantitatively analyze this aspect, we randomly sampled 80 questions from the Food category and evaluated Llava-v1.6-13B’s performance in a generative setting. The model was prompted with the instruction: "This is a question about Korean culture. Answer the given questions briefly and concisely." without access to the original multiple-choice options. Two native Korean evaluators assessed the generated responses using three categories: (1) Correct: The model’s response fully contained the original answer or was accurate to the question, (2) Hallucinated: The model provided incorrect information, and (3) Abstained: The model’s response did not include a direct answer to the question but provided a relevant description of the image and did not attempt to generate an answer hastily (e.g., "The specific type of food and its cultural context would determine when it is commonly served").
The human analysis results are shown in Table 6. We observed that the VLM’s accuracy in the generative setup was relatively lower than in the multiple-choice setup (from 45.07% to 36.25%). This decrease in accuracy is likely because the task is inherently more difficult without predefined options. Interestingly, a large proportion of responses (21.25%) were classified as abstained, indicating that the VLM recognized the need for additional information to answer confidently and refrained from providing a definitive but potentially incorrect response. This behavior suggests that providing the necessary external knowledge could enhance performance. Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
@@ 5.2 Augmenting VLMs with External Knowledge
As shown in Table 7, providing retrieved documents can enhance model performance in the food category, as demonstrated by improved scores in the Retrieved setting compared to None. In cases where proprietary models did not benefit from certain retrieved documents, their performance still surpassed the baseline when given carefully curated Oracle documents. This suggests that retrieval-based augmentation has the potential to strengthen cultural knowledge in VQA tasks, provided that the quality and relevance of the selected documents are refined. We provide the details in the Appendix E.
@@
Table 7: Retrieval-augmented generation results on Food category. Oracle documents were manually selected by the authors.
@@ 6 Conclusion
In this work, we proposed a semi-automated framework for constructing culturally aware benchmarks for vision-language models through human-VLM collaboration, focusing on visual recognition and cultural knowledge application in multi-choice QA. We demonstrated its effectiveness by creating K-Viscuit, revealing a performance gap between proprietary and open-source VLMs in identifying Korean cultural concepts. Through detailed analyses and knowledge augmentation experiments, we specified the impact of cultural understanding on VQA performance and extended our investigation to open-ended answer generation, discussing its limitations. This work highlights the importance of cultural diversity in model evaluation and paves the way for more inclusive VLMs.
@@
Our current framework requires a manual selection of images that match specified concepts, preventing fully automated dataset generation. These human efforts can be alleviated by multimodal retrieval modules to some extent. To this end, it should be predetermined whether current multimodal encoders can sufficiently understand culturally nuanced images and texts. Enhancing retrieval models to better understand and match cultural contexts remains our exciting future work. The manual verification of automatically generated questions also can be a considerable burden. Developing a quality estimation module for generated questions could assist in this process by reducing the workload on human annotators. Additionally, LLMs are highly sensitive to the order of answer options in multiple-choice QA tasks, as discussed in Zheng et al. (2023); Pezeshkpour and Hruschka (2024). While we mitigate positional bias in model evaluation by shuffling the answer options, this approach may not be entirely sufficient. As shown in A.1, VLMs often fail to produce consistent predictions when the positions of the answer options are altered. Future work could explore calibration techniques, such as contrastive decoding or uncertainty-aware selection strategies, to further enhance robustness in multiple-choice settings.
@@
Acknowledgements
This work was supported by a grant of the KAIST-KT joint research project through AI2X Laboratory, Tech Innovation Group , funded by KT (No. D23000019, Research on Multilingual Multicultural Vision-Language Representation Learning) and Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST) and RS-2025-02304967, AI Star Fellowship(KAIST).
Appendix A Additional Analysis
A.1 Evaluating Sensitivity to Option Order in MCQA with Circular Evaluation
Recent studies have shown that many vision-language models (VLMs) exhibit sensitivity to the order of answer choices in multiple-choice QA, where performance varies depending on how options are presented. To investigate this phenomenon, we apply Circular Evaluation, proposed by Liu et al. (2024b). This method systematically shifts the positions of four answer choices and forwards them through the model four times, considering a prediction correct only if all variations yield the same answer.
Figure 9: Order-Sensitivity Analysis. Accuracy comparison of selected models under vanilla and circular evaluations (Liu et al., 2024b). The numbers indicate the absolute and relative performance drops.
Fig 9 presents the accuracy of selected models under both vanilla and circular evaluations. A significant drop in performance is observed across all models when applying circular evaluation, confirming that answer order affects predictions. The relative accuracy decrease ranges from 16.9% to 30.4%, suggesting that models rely on positional cues rather than robust reasoning. Among the evaluated models, Molmo-7B-D exhibits the largest performance degradation, indicating heightened sensitivity to ordering effects. One potential explanation for this behavior is positional bias, where models develop a preference for certain answer positions. These results highlight sensitivity to answer order as a critical challenge in multi-choice QA for vision-language models. Future work should explore mitigation strategies, such as answer re-ranking or training objectives that explicitly reduce order sensitivity.
A.2 Comparison of K-Viscuit with CVQA dataset
Upon reviewing the dataset split, we found that 14 answers (4.8%) were simply "Korea," which may limit the flexibility of different prompt settings for evaluation (e.g., "This is an image taken in Korea"), as "Korea" becomes a shortcut response. Additionally, CVQA occasionally includes distractors that are inconsistent, such as listing (a) "Gwangju," (b) "Seoul," (c) "Seattle," and (d) "Busan," where (c) "Seattle" is an implausible option. This underscores the challenge of ensuring consistency when manually constructing questions and answer choices. While CVQA is a valuable concurrent effort, our work focuses on a systematic approach that leverages VLM collaboration to efficiently build culturally specific datasets.
Category w/o Shot w/ Shot
Food 72.84 77.78
Beverage 76.32 76.32
Game 54.05 56.76
Celebrations 70.83 83.33
Religion 70.59 70.59
Tool 75.56 84.44
Clothes 81.82 86.36
Heritage 69.44 91.67
Architecture 82.50 87.50
Agriculture 67.50 85.00
Total 74.31 81.30
Table 8: Type 2 accuracy with and without one-shot prompting using corresponding Type 1 QA pairs. gpt-4o-mini is used for the experiments.
A.3 Impact of Type 1 Grounding on Type 2 Reasoning
In Table 2, we observe that models tend to perform better on Type 2 questions than on Type 1. This may seem counterintuitive, as Type 1 questions are grounded directly in visual cues. To examine whether grounding in Type 1 content can support Type 2 reasoning, we conducted an experiment where each Type 2 prompt was preceded by its corresponding Type 1 QA pair as a one-shot example. As shown in Table 8, this setup led to consistent gains across categories, improving overall accuracy from 74.31% to 81.30%. These results suggest visual grounding can meaningfully support higher-level cultural reasoning when leveraged appropriately.
Core Concept IDS Chapter No.
Food Food and Beverages (5)
Beverage Food and Beverages (5)
Game Motion (10)
Celebrations Time (14), Religion and belief (22)
Religion Religion and Belief (19)
Tool Basic Actions and Technology (9)
Clothes Clothing and Grooming (6)
Heritage Cognition (17)
Architecture The House (7)
Agriculture Agriculture and Vegetation (8)
Table 9: Mapping between core concepts and IDS chapters. Some concepts share the same chapter due to overlapping semantic coverage.
Appendix B Dataset Construction Details
B.1 Mapping Core Concepts to IDS Chapters
As described in Section 3.1, following Liu et al. (2021), we define the concepts of our dataset based on the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015). Table 9 shows how our 10 core concepts map to chapters in the Intercontinental Dictionary Series (IDS). While some mappings are direct, others reflect one-to-many relationships based on semantic grouping and visual realizability (e.g., Food and Beverage both derive from Chapter 5). This mapping served as the basis for selecting culturally salient and visually grounded concepts in our dataset.
B.2 Guidelines to GPT-4-Turbo for Dataset Annotation
@@
Table 11: A prompt of GPT-4-Turbo used in the annotation of Type 2 questions (i.e., cultural knowledge application) in the Architecture category.
Appendix C Dataset Analyses Details
C.1 Required Knowledge Analysis
@@
C.2 Human Evaluation Details
@@
Appendix D VLM Evaluation Details
@@
Appendix E Retrieval Methodology Details